OpenStudy (anonymous):

Which of the points is a solution to -2x + y = 4? A. (-2, 4) B. (0, 4) C. (2, -2) D. (2, 0)

OpenStudy (anonymous):

We helped you once, time to show some efforts

OpenStudy (anonymous):

studying for test but some problems I have no idea what they are talking about

OpenStudy (amistre64):

same thing as last time; see what fits

OpenStudy (amistre64):

which ordered pair of (x=, y=) makes this equation true?

OpenStudy (anonymous):

A

OpenStudy (amistre64):

can you prove that it would be "a" ? making that choice is one thing, but proving it is another.

OpenStudy (amistre64):

-2x + y = 4 A. (-2, 4) ; so, when (x=-2, y=4) is this equation true? -2(-2) + 4 =? 4 yes or no?

OpenStudy (anonymous):

I think it is yes I just had my paper that i solved the problem on and now I cant find it

OpenStudy (amistre64):

then resolve the problem :) its ok to do things twice

OpenStudy (amistre64):

-2(-2) + 4 = 4 how would we solve this?

OpenStudy (amistre64):

solve is a bad word for that; how would we determine the truth of this?

OpenStudy (amistre64):

-2 * -2 = ?

OpenStudy (anonymous):

4

OpenStudy (amistre64):

good, -2(-2) + 4 = 4 4 +4 = 4 what is 4+4 equal?

OpenStudy (anonymous):

8

OpenStudy (amistre64):

ok -2(-2) + 4 = 4 4 +4 = 4 8 = 4 does 8 equal 4 make any sense? is it a true statement?

OpenStudy (anonymous):

no

OpenStudy (amistre64):

then since A makes a false statement, and lies to us; we say that A is NOT a solution

OpenStudy (amistre64):

which of the other 3 options will NOT lie to us?

OpenStudy (anonymous):

b 0,4 because it doesnt make a false statement

OpenStudy (amistre64):

i would agree with that, and I have proof :) 2x + y = 4 0 4 ----------- 0 + 4 = 4 ; IS a true statement, therefore this IS a solution